Question # 02

A 90-kg football player stands on a frictionless lake of ice. He catches a 3-kg football and then moves at 400 cm/s. What was the initial velocity of the football? 


1
Expert's answer
2021-04-05T11:11:01-0400
"m_bv_b+m_pv_p=(m_b+m_p)v_c,""m_bv_b+0=(m_b+m_p)v_c,""v_b=\dfrac{(m_b+m_p)v_c}{m_b}=\dfrac{(3\ kg+90\ kg)\cdot4\ \dfrac{m}{s}}{3\ kg}=124\ \dfrac{m}{s}."